# Stability Analysis of Cosmological Perturbations in the Bumblebee Model: Parameter Constraints and Gravitational Waves

### v2

@@ Appendix A The specific forms of some quantities
$Q_{\phi_{h}}$ $=$ $\frac{3}{2}\xi b_{t}^{2}a^{3}\ddot{E}+3a^{3}H\dot{E}-a\partial^{2}E-3(1-\xi b_{t}^{2})a^{3}\dot{H}E+\frac{\xi b_{t}^{2}}{2}a^{3}\partial^{2}\ddot{\alpha}+a^{3}H\partial^{2}\dot{\alpha}-(1-\xi b_{t}^{2})a^{3}\dot{H}\partial^{2}\alpha$ (147)
$-\xi b_{t}^{2}a\partial^{2}\dot{\varphi}_{h}-2(1-\xi b_{t}^{2})aH\partial^{2}\varphi_{h}-3\xi b_{t}a^{3}H\dot{\phi}_{b}+\xi b_{t}a\partial^{2}\phi_{b}+\big(-3\xi b_{t}a^{3}(H^{2}-\dot{H})+4\kappa b_{t}^{3}a^{3}\overline{V}_{,b^{2}b^{2}}\big)\phi_{b}$
$-2\xi b_{t}^{2}a\partial^{2}\phi_{h}+\big(-6H^{2}+3\xi b_{t}^{2}(H^{2}-3\dot{H})-4\kappa b_{t}^{4}\overline{V}_{,b^{2}b^{2}}\big)a^{3}\phi_{h}+2\xi b_{t}aH\partial^{2}\varphi_{b}-\kappa\overline{\rho}_{m,n}\phi_{m},$
$Q_{\varphi_{h}}$ $=$ $\partial^{2}\Big(\xi b_{t}^{2}\dot{\phi}_{h}-(2-3\xi b_{t}^{2})H\phi_{h}+(1-\xi b_{t}^{2})\dot{E}-\xi b_{t}\dot{\phi}_{b}-\xi b_{t}H\phi_{b}+2(1-\xi b_{t}^{2})\dot{H}\varphi_{h}-\frac{\kappa\overline{\rho}_{m,n}}{a^{3}}\varphi_{m}\Big),$ (148)
$Q_{E}$ $=$ $3\xi b_{t}^{2}\ddot{\phi}_{h}-6(1-3\xi b_{t}^{2})H\dot{\phi}_{h}-\frac{2}{a^{2}}\partial^{2}\phi_{h}-\big(9(2-3\xi b_{t}^{2})H^{2}+3(4-5\xi b_{t}^{2})\dot{H}\big)\phi_{h}+3(1-\xi b_{t}^{2})\ddot{E}$ (149)
$+9(1-\xi b_{t}^{2})H\dot{E}-\frac{1}{a^{2}}\partial^{2}E-\frac{9\kappa\overline{J}^{2}\overline{\rho}_{m,nn}}{2a^{6}}E+\partial^{2}\Big((1-\xi b_{t}^{2})\ddot{\alpha}+3(1-\xi b_{t}^{2})H\dot{\alpha}-\frac{3\kappa\overline{J}^{2}\overline{\rho}_{m,nn}}{2a^{6}}\alpha\Big)$
$-3\xi b_{t}\big(\ddot{\phi}_{b}+4H\dot{\phi}_{b}+(3H^{2}+\dot{H})\phi_{b}\big)-\frac{2}{a^{2}}\partial^{2}\big((1-\xi b_{t}^{2})\dot{\varphi}_{h}+(1-\xi b_{t}^{2})H\varphi_{h}-\xi b_{t}\dot{\varphi}_{b}-\xi b_{t}H\varphi_{b}\big)$
$+\frac{3\kappa\overline{J}\overline{\rho}_{m,nn}}{a^{6}}\phi_{m},$
$Q_{\alpha}$ $=$ $\partial^{2}\Big(\xi b_{t}^{2}\ddot{\phi}_{h}-2(1-3\xi b_{t}^{2})H\dot{\phi}_{h}-\big(3(2-3\xi b_{t}^{2})H^{2}+(4-5\xi b_{t}^{2})\dot{H}\big)\phi_{h}+(1-\xi b_{t}^{2})(\ddot{E}+3H\dot{E})$ (150)
$\qquad-\frac{3\kappa\overline{J}^{2}\overline{\rho}_{m,nn}}{2a^{6}}E-\xi b_{t}\big(\ddot{\phi}_{b}+4H\dot{\phi}_{b}+(3H^{2}+\dot{H})\phi_{b}\big)+\frac{\kappa\overline{J}\overline{\rho}_{m,nn}}{a^{6}}\phi_{m}-\frac{\kappa\overline{J}^{2}\overline{\rho}_{m,nn}}{2a^{6}}\partial^{2}\alpha\Big),$
$Q_{\phi_{b}}$ $=$ $-3\xi b_{t}(\ddot{E}+2H\dot{E})-\xi b_{t}\partial^{2}(\ddot{\alpha}+2H\dot{\alpha})+6\xi b_{t}H\dot{\phi}_{h}+\frac{2\xi b_{t}}{a^{2}}\partial^{2}\phi_{h}+4b_{t}\big(3\xi(H^{2}+\dot{H})+2\kappa b_{t}^{2}\overline{V}_{b^{2}b^{2}}\big)\phi_{h}$ (151)
$+\frac{2\xi b_{t}}{a^{2}}\partial^{2}\dot{\varphi}_{h}+\frac{2\kappa}{a^{2}}\partial^{2}\dot{\varphi}_{b}-\frac{2\kappa}{a^{2}}\partial^{2}\phi_{b}-8\kappa b_{t}^{2}\overline{V}_{,b^{2}b^{2}}\phi_{b},$
@@
The specific forms of the coefficients in the constraint equation (97) are given by:
$F_{1}(k,6)$ $=$ $\xi b_{t}^{4}(\xi-2\kappa)ak^{6}+b_{t}^{2}a^{3}\big(3(2\kappa-4\xi-\kappa\xi b_{t}^{2}+4\xi^{2}b_{t}^{2})H^{2}+3\xi b_{t}^{2}(3\kappa-2\xi)\dot{H}+4\kappa^{2}b_{t}^{4}\overline{V}_{,b^{2}b^{2}}\big)k^{4}$ (156)
$+3a^{5}\big(6(2-3\xi b_{t}^{2}+\xi^{2}b_{t}^{4})H^{4}+3\xi^{2}b_{t}^{4}\dot{H}^{2}+2b_{t}^{2}(1-\xi b_{t}^{2})(15\xi\dot{H}+4\kappa b_{t}^{2}\overline{V}_{,b^{2}b^{2}})H^{2}\big)k^{2}$
$-18(1-\xi b_{t}^{2})\big(3(2-\xi b_{t}^{2})H^{2}+9\xi b_{t}^{2}\dot{H}+4\kappa b_{t}^{4}\overline{V}_{,b^{2}b^{2}}\big)a^{7}H^{2}\dot{H},$
$F_{2}(k,6)$ $=$ $\kappa b_{t}^{2}a\big(-2\xi b_{t}^{2}k^{2}+a^{2}(3(2-\xi b_{t}^{2})H^{2}+9\xi b_{t}^{2}\dot{H}+4\kappa b_{t}^{4}\overline{V}_{,b^{2}b^{2}})\big)k^{4},$ (157)
$F_{3}(k,6)$ $=$ $2\xi b_{t}^{2}aH\big(-\xi b_{t}^{2}k^{2}+a^{2}(3(2-\xi b_{t}^{2})H^{2}+6\xi b_{t}^{2}\dot{H}+4\kappa b_{t}^{4}\overline{V}_{,b^{2}b^{2}})\big)k^{4},$ (158)
$F_{4}(k,6)$ $=$ $-\xi^{2}b_{t}^{5}a(k^{2}-3a^{2}\dot{H})k^{4},$ (159)
$F_{5}(k,6)$ $=$ $-2b_{t}(1-\xi b_{t}^{2})(k^{2}-3a^{2}\dot{H})\big(-\xi b_{t}^{2}k^{2}+a^{2}(3(2-\xi b_{t}^{2})H^{2}+9\xi b_{t}^{2}\dot{H}+4\kappa b_{t}^{4}\overline{V}_{,b^{2}b^{2}})\big)aHk^{2},$ (160)
$F_{6}(k,2)$ $=$ $-3\kappa b_{t}\big(-2\xi b_{t}^{2}k^{2}+a^{2}(3(2-\xi b_{t}^{2})H^{2}+9\xi b_{t}^{2}\dot{H}+4\kappa b_{t}^{4}\overline{V}_{,b^{2}b^{2}})\big)a^{2}H\overline{\rho}_{m,n},$ (161)
$F_{7}(k,4)$ $=$ $\kappa b_{t}\big(-\xi b_{t}^{2}k^{2}+a^{2}(3(2-\xi b_{t}^{2})H^{2}+6\xi b_{t}^{2}\dot{H}+4\kappa b_{t}^{4}\overline{V}_{,b^{2}b^{2}})\big)\overline{\rho}_{m,n}k^{2}.$ (162)

### v3

@@ Appendix A The specific forms of some quantities
$Q_{\phi_{h}}$ $=$ $\frac{3}{2}(\xi+2\sigma)b_{t}^{2}a^{3}\ddot{E}+3(1+3\sigma b_{t}^{2})a^{3}H\dot{E}-(1+\sigma b_{t}^{2})a\partial^{2}E-3(1-(\xi+\sigma)b_{t}^{2})a^{3}\dot{H}E+\frac{(\xi+2\sigma)b_{t}^{2}}{2}a^{3}\partial^{2}\ddot{\alpha}$ (147)
$+(1+3\sigma b_{t}^{2})a^{3}H\partial^{2}\dot{\alpha}-(1-(\xi+\sigma)b_{t}^{2})a^{3}\dot{H}\partial^{2}\alpha-(\xi+2\sigma)b_{t}^{2}a\partial^{2}\dot{\varphi}_{h}-2(1-(\xi-\sigma)b_{t}^{2})aH\partial^{2}\varphi_{h}$
$-3(\xi+2\sigma)b_{t}a^{3}H\dot{\phi}_{b}+(\xi+2\sigma)b_{t}a\partial^{2}\phi_{b}+\big(-3\xi(H^{2}-\dot{H})+6\sigma(H^{2}+\dot{H})+4\kappa b_{t}^{2}\overline{V}_{,b^{2}b^{2}}\big)b_{t}a^{3}\phi_{b}$
$-2(\xi+2\sigma)b_{t}^{2}a\partial^{2}\phi_{h}-\Big(3\big(2-(\xi-8\sigma)b_{t}^{2}\big)H^{2}+9(\xi+2\sigma)b_{t}^{2}\dot{H}+4\kappa b_{t}^{4}\overline{V}_{,b^{2}b^{2}}\Big)a^{3}\phi_{h}$
$+2\xi b_{t}aH\partial^{2}\varphi_{b}-\kappa\overline{\rho}_{m,n}\phi_{m},$
$Q_{\varphi_{h}}$ $=$ $\partial^{2}\Big((\xi+2\sigma)b_{t}^{2}\dot{\phi}_{h}-(2-3\xi b_{t}^{2})H\phi_{h}+(1-(\xi+\sigma)b_{t}^{2})\dot{E}-(\xi+2\sigma)b_{t}\dot{\phi}_{b}-(\xi-2\sigma)b_{t}H\phi_{b}$ (148)
$+2(1-(\xi+\sigma)b_{t}^{2})\dot{H}\varphi_{h}-\frac{\kappa\overline{\rho}_{m,n}}{a^{3}}\varphi_{m}\Big),$
$Q_{E}$ $=$ $3(\xi+2\sigma)b_{t}^{2}\ddot{\phi}_{h}-6\big(1-3(\xi+\sigma)b_{t}^{2}\big)H\dot{\phi}_{h}-\frac{2(1+\sigma b_{t}^{2})}{a^{2}}\partial^{2}\phi_{h}-3\Big(3(2-3\xi b_{t}^{2})H^{2}+(4-(5\xi+2\sigma)b_{t}^{2})\dot{H}\Big)\phi_{h}$ (149)
$+3(1-(\xi+\sigma)b_{t}^{2})\ddot{E}+9(1-(\xi+\sigma)b_{t}^{2})H\dot{E}-\frac{(1-\sigma b_{t}^{2})}{a^{2}}\partial^{2}E-\frac{9\kappa\overline{J}^{2}\overline{\rho}_{m,nn}}{2a^{6}}E+\frac{3\kappa\overline{J}\overline{\rho}_{m,nn}}{a^{6}}\phi_{m}$
$+\partial^{2}\Big((1-(\xi+\sigma)b_{t}^{2})\ddot{\alpha}+3(1-(\xi+\sigma)b_{t}^{2})H\dot{\alpha}-\frac{3\kappa\overline{J}^{2}\overline{\rho}_{m,nn}}{2a^{6}}\alpha\Big)-3(\xi+2\sigma)b_{t}\ddot{\phi}_{b}-12(\xi+\sigma)b_{t}H\dot{\phi}_{b}$
$-3(\xi-2\sigma)b_{t}(3H^{2}+\dot{H})\phi_{b}+\frac{4\sigma b_{t}}{a^{2}}\partial^{2}\phi_{b}-\frac{2}{a^{2}}\partial^{2}\big((1-(\xi+\sigma)b_{t}^{2})\dot{\varphi}_{h}+(1-(\xi+\sigma)b_{t}^{2})H\varphi_{h}\big)$
$+\frac{2\xi b_{t}}{a^{2}}\partial^{2}\big(\dot{\varphi}_{b}+H\varphi_{b}\big),$
$Q_{\alpha}$ $=$ $\partial^{2}\Big((\xi+2\sigma)b_{t}^{2}\ddot{\phi}_{h}-2(1-3(\xi+\sigma)b_{t}^{2})H\dot{\phi}_{h}-\big(3(2-3\xi b_{t}^{2})H^{2}+(4-(5\xi+2\sigma)b_{t}^{2})\dot{H}\big)\phi_{h}-\frac{3\kappa\overline{J}^{2}\overline{\rho}_{m,nn}}{2a^{6}}E$ (150)
$+\big(1-(\xi+\sigma)b_{t}^{2}\big)(\ddot{E}+3H\dot{E})-(\xi+2\sigma)b_{t}\ddot{\phi}_{b}-4(\xi+\sigma)b_{t}H\dot{\phi}_{b}-(\xi-2\sigma)b_{t}(3H^{2}+\dot{H})\phi_{b}$
$+\frac{\kappa\overline{J}\overline{\rho}_{m,nn}}{a^{6}}\phi_{m}-\frac{\kappa\overline{J}^{2}\overline{\rho}_{m,nn}}{2a^{6}}\partial^{2}\alpha\Big),$
$Q_{\phi_{b}}$ $=$ $-3(\xi+2\sigma)b_{t}\ddot{E}-6(\xi+4\sigma)b_{t}H\dot{E}+\frac{4\sigma b_{t}}{a^{2}}\partial^{2}E-(\xi+2\sigma)b_{t}\partial^{2}\ddot{\alpha}-2(\xi+4\sigma)b_{t}H\partial^{2}\dot{\alpha}+6(\xi+2\sigma)b_{t}H\dot{\phi}_{h}$ (151)
$+\frac{2(\xi+2\sigma)b_{t}}{a^{2}}\partial^{2}\phi_{h}+4b_{t}\big(3(\xi+4\sigma)H^{2}+3(\xi+2\sigma)\dot{H}+2\kappa b_{t}^{2}\overline{V}_{b^{2}b^{2}}\big)\phi_{h}+\frac{2(\xi+2\sigma)b_{t}}{a^{2}}\partial^{2}\dot{\varphi}_{h}+\frac{8\sigma b_{t}}{a^{2}}H\partial^{2}\varphi_{h}$
$+\frac{2\kappa}{a^{2}}\partial^{2}\dot{\varphi}_{b}-\frac{2\kappa}{a^{2}}\partial^{2}\phi_{b}-8\kappa b_{t}^{2}\overline{V}_{,b^{2}b^{2}}\phi_{b},$